In an experiment on a venturi meter installed in a 300 mm diameter pipe, the mercury in the differential U-tube manometer connected to the venturi meter deflects 73 mm when 73 l/s of water is flowing in the pipe. If the diameter of the throat of the venturi is 150 mm, calculate Cd for the venturi


1
Expert's answer
2021-11-19T00:07:28-0500

d1= 0.3m

d2=0.15m

x=0.073m

Q=0.073m^3/sec

To calculate:

Cd=?

"a_1=\pi /4 (0.3)^2=0.0706m^2
\space ,a_2 \pi /4 (0.15)^2=0.0176m^2"

"h=x(S_g\/S_w-1)=0.073(13.6\/1 -1)=0.9198"


"Q=C_d \frac{a_1 a_2(2gh)^{1/2}}{(a_1^2-a_2^2)^{1/2}}"


"C_d=\frac{0.073}{0.0773}"


"C_d=0.943 \space \text{coeficiant of discharge}"